# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ 4.3.1   LR Regime
$\mathscr{G}_{\text{LR}}=\beta\sum_{a=2}^{n}\frac{\Omega_{a}^{2}}{-\lambda_{a}}\,,$ (S120)
with $\Omega_{a}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}$, where the vectors $\boldsymbol{m}^{\alpha}$ correspond to the eigensystem discussed in discussed in Sec. 4.1, i.e. in (S107). On the other hand, here $\boldsymbol{v}=\frac{\beta\theta}{2\sqrt{n}}(1,0,\ldots,0)$. The latter is obtained by construction using (S66), (S71) and (S72). Thus,
@@
$\sum_{a=2}^{n/2}\frac{1}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}\simeq\frac{n^{2}}{12}+\mathcal{O}(1)\,,$ (S121)
@@
$\mathscr{G}_{\text{LR}}\simeq\frac{\beta}{12k}\left(\frac{\theta}{2}\right)^{2}\,.$ (S122)
If we wanted to approximate the result above for a large odd $n$, we would need to add a term proportional to $1/4n^{2}$, which is of second order, hence the leading order is still captured by expression (S122).
We can also study the optimal distribution using (S77),
$\Delta\bm{p}^{*}_{\mathrm{LR}}=\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}={\beta}\sum_{a=2}^{n}\frac{\Omega_{a}^{\dagger}D\boldsymbol{m}^{\alpha}}{-\lambda_{a}}=\frac{\beta\theta}{4kn^{2}}\sum_{a=2}^{n}\frac{\boldsymbol{u}^{\alpha}}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}$ (S123)
@@
Figure S14: The optimal $\Delta\bm{p}^{*}_{\mathrm{LR}}$ (the deviation of the optimal distribution from the uniform baseline steady state) from (S123). Note that there optimal distribution does not lead to a cyclic current.
Expression (S79) suggests sufficiency conditions for when the optimum will be in the LR regime. Note that, in this case, $\max_{a>1}\Omega_{a}={\theta}/{2n}$ while
@@
Then, using $\|D\|=1/\sqrt{n}$, condition (S79) will read as:
$\frac{\theta}{2n}\ll\frac{4k\sqrt{n}}{\beta(n-1)}\sin^{2}\left(\frac{\pi}{n}\right)\Leftrightarrow\theta\ll\frac{8\sqrt{n}k}{\beta}\left(\frac{n}{n-1}\right)\sin^{2}\left(\frac{\pi}{n}\right)\,.$ (S124)
@@
$\theta\ll\frac{8\pi^{2}}{n^{3/2}}\frac{k}{\beta}\,.$ (S125)
Expression (S79), however, is not necessarily a tight bound. That is, the result hereby obtained is a sufficient condition for the LR regime to be valid, but not a necessary one.
4.3.2   D and ND Regimes
The optimal mesostate is $i^{*}=1$. Hence $\mathscr{G}_{\text{D}}=\phi_{1}=\theta$ and $(\bm{p}^{*}_{\mathrm{D}})_{i}=\delta_{i1}$. Next, by following a similar procedure as above, we denote $\Delta\phi_{2}=\Delta\phi_{n}=\theta$, $R^{b}_{21}=R^{b}_{n1}=k$, $R^{b}_{11}=-2k$ and $R^{b}_{j1}=0$ for all $j\neq 2,n$. Thus, for $i\neq i^{*}$, we obtain:
@@ 4.3.2   D and ND Regimes
Note that, for this example, ${\bm{\pi}^{b}}^{T}{\phi^{M}}={\theta}/{n}$. Therefore,
@@
$=\theta-2k{\beta}^{-1}\left[1+\ln\left(\beta\theta k^{-1}-2\right)\right]\,.$ (S127)
Quadratic optimization lemma
@@
*$\mathscr{L}^{*}=\max_{\boldsymbol{z}\in\mathbb{R}^{n}:\boldsymbol{u}^{T}\boldsymbol{z}=0}\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v},$ (S128)
@@
$\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v}=\boldsymbol{z}^{T}M\boldsymbol{z}+\boldsymbol{v}^{T}M^{+}M\boldsymbol{z}+\boldsymbol{z}^{T}MM^{+}\boldsymbol{v}=\left(\boldsymbol{z}+M^{+}\boldsymbol{v}\right)^{T}M\left(\boldsymbol{z}+M^{+}\boldsymbol{v}\right)-\boldsymbol{v}^{T}M^{+}\boldsymbol{v}$ (S130)

### 12 Mar 2024 v4

@@ IV.3.1 LR Regime
$\mathscr{G}_{\text{LR}}=\beta\sum_{a=2}^{n}\frac{\Omega_{a}^{2}}{-\lambda_{a}}\,,$ (S4.119)
with $\Omega_{a}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}$, where the vectors $\boldsymbol{m}^{\alpha}$ correspond to the eigensystem discussed in discussed in Sec. IV.1, i.e. in (S4.106). On the other hand, here $\boldsymbol{v}=\frac{\beta\theta}{2\sqrt{n}}(1,0,\ldots,0)$. The latter is obtained by construction using (S3.65), (S3.70) and (S3.71). Thus,
@@
$\sum_{a=2}^{n/2}\frac{1}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}\simeq\frac{n^{2}}{12}+\mathcal{O}(1)\,,$ (S4.120)
@@
$\mathscr{G}_{\text{LR}}\simeq\frac{\beta}{12k}\left(\frac{\theta}{2}\right)^{2}\,.$ (S4.121)
If we wanted to approximate the result above for a large odd $n$, we would need to add a term proportional to $1/4n^{2}$, which is of second order, hence the leading order is still captured by expression (S4.121).
We can also study the optimal distribution using (S3.76),
$\Delta\bm{p}^{*}_{\mathrm{LR}}=\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}={\beta}\sum_{a=2}^{n}\frac{\Omega_{a}^{\dagger}D\boldsymbol{m}^{\alpha}}{-\lambda_{a}}=\frac{\beta\theta}{4kn^{2}}\sum_{a=2}^{n}\frac{\boldsymbol{u}^{\alpha}}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}$ (S4.122)
@@
Figure S16: The optimal $\Delta\bm{p}^{*}_{\mathrm{LR}}$ (the deviation of the optimal distribution from the uniform baseline steady state) from (S4.122). Note that there optimal distribution does not lead to a cyclic current.
Expression (S3.78) suggests sufficiency conditions for when the optimum will be in the LR regime. Note that, in this case, $\max_{a>1}\Omega_{a}={\theta}/{2n}$ while
@@
Then, using $\|D\|=1/\sqrt{n}$, condition (S3.78) will read as:
$\frac{\theta}{2n}\ll\frac{4k\sqrt{n}}{\beta(n-1)}\sin^{2}\left(\frac{\pi}{n}\right)\Leftrightarrow\theta\ll\frac{8\sqrt{n}k}{\beta}\left(\frac{n}{n-1}\right)\sin^{2}\left(\frac{\pi}{n}\right)\,.$ (S4.123)
@@
$\theta\ll\frac{8\pi^{2}}{n^{3/2}}\frac{k}{\beta}\,.$ (S4.124)
Expression (S3.78), however, is not necessarily a tight bound. That is, the result hereby obtained is a sufficient condition for the LR regime to be valid, but not a necessary one.
IV.3.2 D and ND Regimes
The optimal mesostate is $i^{*}=1$. Hence
$\mathscr{G}_{\text{D}}=\phi_{1}=\theta\qquad\qquad(\bm{p}^{*}_{\mathrm{D}})_{i}=\delta_{i1}\,.$
Next, by following a similar procedure as above, we denote $\Delta\phi_{2}=\Delta\phi_{n}=\theta$, $R^{b}_{21}=R^{b}_{n1}=k$, $R^{b}_{11}=-2k$ and $R^{b}_{j1}=0$ for all $j\neq 2,n$. Thus, for $i\neq i^{*}$, we obtain:
@@ IV.3.2 D and ND Regimes
The harvesting rate attainable in this regime can be calculated using Eq. (S3.97),
@@
$=\theta-2k{\beta}^{-1}\left[1+\ln\left(\beta\theta k^{-1}-2\right)\right]\,.$ (S4.126)
V Quadratic optimization lemma
@@
*$\mathscr{L}^{*}=\max_{\boldsymbol{z}\in\mathbb{R}^{n}:\boldsymbol{u}^{T}\boldsymbol{z}=0}\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v},$ (S5.127)
@@
$\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v}=\boldsymbol{z}^{T}M\boldsymbol{z}+\boldsymbol{v}^{T}M^{+}M\boldsymbol{z}+\boldsymbol{z}^{T}MM^{+}\boldsymbol{v}=\left(\boldsymbol{z}+M^{+}\boldsymbol{v}\right)^{T}M\left(\boldsymbol{z}+M^{+}\boldsymbol{v}\right)-\boldsymbol{v}^{T}M^{+}\boldsymbol{v}$ (S5.129)